# Masked Mixers for Language Generation and Retrieval

### v1

@@ 3.7 Autoregressive inference for masked mixers
Cross-entropy loss for all-next-token training (using identical tokenizers) has allowed for a direct comparison of the abilities of masked mixers and transformers on the same dataset. The ultimate goal of causal language model training is often language generation, and now we compare the language generated by transformers and masked mixers.
Masked mixers cannot be inferenced in a manner analogous to transformers because these models have a fixed feedforward context window size, at least without modification to the present implementation. This means that a full forward pass is required for each token generation such that the masked mixer inferences with time complexity $O(n^{3}*d)$ where $n$ is the number of tokens and $d=d_{model}$. In practice masked mixers inference only slightly slower than transformers that use key-value caching and are thus $O(n^{2}*d)$ because of the much lower constant factors for the mixers.
@@
A Python implementation of this approach for is as follows:
@@
Using this inferencing method, we find that masked mixers generate similarly readable outputs to transformers at the same CLM cross-entropy loss value. For example, given the prompt
One day, a little boy named Tim went to play with his friend, Sam. They wanted to play a game with a ball. The game was to see who could get the best score. Tim threw the ball and Sam tried to catch it. Sam missed the ball and it went far away. Sam was angry. He said, "I don’t like this game. I want to lie down and rest. "Tim said, "Okay, let’s lie down and rest. Then we can try again to get a good score." They lay down under a big tree and looked at the sky. They saw birds flying and clouds moving.After a while, they got up and a 1024-dim flat masked mixer with validation cross-entropy loss of 1.81 yields a fairly coherent story completion:
@@
whereas a 256-dim transformer with a validation loss of 1.91 loss yields the the following completion,
@@
with somewhat more deviation from the prompt’s plot.
4 Masked mixers are more effective for retrieval than transformers
@@
On the other hand, matching a sequences of tokens to another sequence of tokens may be thought of as an approximately bijective function which would not be expected to benefit from information reduction or non-invertibility. Masked mixers have been shown to be effectively invertible for small inputs and in that respect retain much more input information than mixers, suggesting that these models are better suited to retrieval than models with attention assuming that retrieval requires most input information. In this section we test this idea experimentally, and find that indeed last-hidden-layer last-token embeddings from masked mixers are far superior for retrieval tasks compared to the same embeddings from transformers.
@@ 4.1 Synthetic dataset generation and embedding
Testing retrieval ability using our existing models entails using the dataset these models were trained on, especially as the tokenizers used for these models will generally be unsuitable for other text corpora. To that end we generated a synthetic summary sentence for the first 200k stories in the TinyStories dataset using Llama-3 (8b) Instruct [5] by prompting that model to read each story and give a very short summary of that story. Due to the relatively simple nature of these stories, we found that this relatively small model was more than capable of accurately and concisely summarizing each story.
Each story and its corresponding summary were fed to either a Llama-style transformer or a masked mixer, and the last hidden layer activations were saved as the model’s embedding of its input. The embedding from the transformer’s last token was used, whereas the second-to-last token embedding was used for the masked mixer. This is because the ’last’ masked mixer token is untrained due to the index shift that occurs during all-next-token training of these models. An example of a TinyStory and its one-sentence summary is found below:
Text: Once upon a time, there was a smelly old tree. By the tree, there was a big hole.
In the hole, there was a shiny coin. A boy named Tim saw the coin and wanted it.
Tim said to his friend, "I want that coin, but the tree is smelly and rot."
His friend said, "Be brave, Tim! You can get the coin." So, Tim went to the smelly tree
[…]
little mouse. The mouse said, "I will help you get the coin, but you must help me too."
Tim agreed, and they got the coin together. Tim was happy, and the mouse was happy too.
And they both learned that even in smelly places, good things can happen.
Summary: A boy named Tim overcomes his fear of a smelly old tree to retrieve a shiny
coin with the help of a friendly mouse.
4.2 Retrieval model architecture and training approach
A standard method for transformer model retrieval is to use the cosine distance on the output, defined as follows:
$d(x,y)=1-\cos(x,y)=1-\frac{x\cdot y}{||x||\;||y||}$ (7)
which is usually implemented by first normalizing $x,y$ such that $||x||=||y||=1$ and then performing matrix matrix multiplication on matrices $X,Y$ of concatenated vectors of $x,y$. Here the
It has been found that cosine distance on an output is in some ways well-aligned with the dot product self-attention transformations (c.f. [14]) and thus this metric would not be expected to lead to as accurate retrieval for attentionless models like mixers. Indeed this is found to be the case, and preliminary experimentation revealed that $L^{1}$ and $L^{2}$ metrics are also unsuitable such that we instead decided to directly train for retrieval using fixed embeddings from previously-trained masked mixers and transformers.
Training models for the purpose of retrieval is not a new idea, with most state-of-the-art retrieval models receiving training to minimize the cosine distance for matching pairs of sequences while maximizing the distance for non-matching pairs. This training typically proceeds by modifying the parameters of the embedding model directly [25] which in some sense is very inefficient because one forward pass is required for each input for each batch of matching and non-matching sequences.
One can instead train far more efficiently by effectively compressing each input using a trained generative model, and then training a separate model on these embeddings such that a single forward pass is required for comparison of all inputs in each batch of embeddings of size $c$. This method has the added bonus of allowing one to examining the effectiveness of training one type of model on the inputs from various architectures without having to account for differences in these architectures.
@@ 4.2 Retrieval model architecture and training approach
For our retrieval model we chose to implement a mixer with bidirectional convolutions rather than masked ones, similar to MLP mixers but with only one layer rather than two between sequence elements. This was chosen because a comparison of all possible matches requires information from all input elements by definition, making transformer models unsuitable. This was experimentally verified, as transformers were found to be very poor retrieval models even for small comparison contexts (Table S8). Word -token embedding layers are removed (as the inputs are embeddings already) as are language heads, and instead each last hidden layer has a $d_{model}\to 1$ transformation where the unitary output corresponds to the likelihood (after Softmax transformation) of that input embedding being the correct match. Softmax transformation serves to stabilize the retrieval model’s loss, as increasing all logit values does not decrease the cross-entropy loss. We make use of the Pytorch Cross-Entropy loss function’s compatibility with class probabilities for this model.
@@
The outputs after Softmax transformation are compared to the true distribution via standard cross-entropy loss, which is then back-propagated through the retrieval model but not further to the generative model or input embedding. Note that the logits are not Softmax-transformed during inference for speed, as this operation has no effect on topk choice without noise introduction. We use a relatively modest batch size compared to most retrieval model training approaches (512 total). See Figure 10 for a representation of the retrieval model.
Figure 10: Left: Retrieval model architecture. Upper Right: Learning curves for the retrieval model with embeddings from CLM-trained masked mixers or transformers. Lower Right: Top-1 accuracy for evaluation samples (20k)
4.3 Embeddings from masked mixers result in much more accurate retrieval compared to embeddings from transformers
In contrast to the rest of this work, the efficiency of an embedding for retrieval modeling is not measured in loss per compute amount or even loss per number of samples seen but instead minimum evaluation cross-entropy loss over many (200) epochs of training. Retrieval model training was found to have an unusual and stereotypically Sigmoid loss over time. In the worst case the retrieval model fails to break symmetry and predicts one identical output for all embedding inputs, and this corresponds to little decrease in loss and may coincide with gradient explosion and overflow after the first hundred or so epochs of training. More commonly, the retrieval model will train to near-perfect accuracy on training data and the generalization to the evaluation dataset may be examined (Figure 10).
Embeddings from masked mixers after causal language model training lead to far lower evaluation set cross-entropy loss minimization than embeddings from transformers with the same training approach (Figure 10, Table S9, and Table S10). This corresponds to far higher top-1 retrieval performance on the evaluation dataset 10. In particular, the best transformer embedding observed for 128-context retrieval has a cross-entropy loss of 2.28 and top-1 accuracy of 43.2%, compared to the best masked mixer embedding with a loss of 0.87 and top-1 accuracy of 85.8%.
Larger datasets predictably lead to increased accuracy for masked mixer embeddings: increasing the training dataset size by a factor of around 3 leads to a drop in 128-context comparison cross-entropy evaluation loss from 0.23 to 0.10 in the same validation dataset (for the 512-dimensional masked mixer CLM embeddings sampled with replacement). This corresponds to a top-1 accuracy of 97.0% for that context window size, an increase from 85.8% achieved with 200k samples.
4.4 Untrained or non-CLM trained models embeddings yield poor retrieval
We tested the ability of embeddings from untrained masked mixers to be used for effective retrieval modeling, as would be predicted by the more accurate input representation of untrained versus trained masked mixers (Figure 3). If it were indeed the case that one could forego causal language model training for the embedding model a great deal of compute could be saved particularly for larger datasets or instances where custom tokenizers are required. A masked mixer was initialized and then used produced a set of story and summary embeddings without the CLM-style training that normally precedes embedding generation. The standard bidirectional mixer retrieval model used previously trained very poorly when using these embeddings for matching batches of context size 128 or even 32, with practically no decrease in loss observed (Table S9). This suggests that training endows the embedding model with the ability to capture important parts of the language input, and that accurate input representation does not have a simple relationship to retrieval ability.
We also investigated whether the type of training is important to this retrieval process, and in particular whether or not all-next-token style causal language model training is necessary for effective retrieval. This can be done by swapping out the masked mixer generative model for one that learns to represent an input via a non-trivial autoencoding process. We implement this autoencoder using a encoder-decoder model where a standard masked mixer acts as the encoder, the last hidden state of the last token is taken as the embedding, and a decoder composed of a stack of masked mixer modules transforms this embedding (repeated for each input word) into the input sequence as shown in 11.
Figure 11: Bidirectional Mixer Autoencoder architecture for $d_{m}=1024$
This mixer autoencoder reaches a cross-entropy loss on TinyStories comparable to non-optimized transformers and mixers (albeit with around twice the number of training steps required) and very little or no overfitting, indicating that the embedding created by the model’s encoder provides generalizable information on the input to the decoder, and in particular enough information for the decoder to reconstruct the entire input sequence approximately as accurately as trained all-next-token causal language models. It comes as some surprise, therefore, that the embeddings from this autoencoder are no better than embeddings from an untrained mixer when applied to retrieval training, even for the relatively small n=32 context sizes (Table S9).
It should be noted that this retrieval model approach is also efficient for inference as well as training: with batching, a retrieval model with a ‘context’ window of 128 samples with a batch size of 128 requires less than 6 GB vRAM, but is capable of matching a story to $16384128^{2}=$ 16384 samples in a two forward passes.

### v2

@@ 4.2 Autoregressive inference for masked mixers
Masked mixers cannot be inferenced in a manner analogous to transformers because these models have a fixed feedforward context window size, at least without modification (convolutional weight restrictions for example) to the present implementation. This means that a full forward pass is required for each token generation such that the masked mixer inferences with time complexity $O(n^{3}*d)$ where $n$ is the number of tokens and $d=d_{model}$. In practice masked mixers inference only slightly slower than transformers that use key-value caching and are thus $O(n^{2}*d)$ because of the much lower constant factors for the mixers.
@@
A Python implementation of this approach is as follows:
@@
Using this inferencing method we find that masked mixers generate similarly readable outputs to TinyStories prompts as transformers at the same validation cross-entropy loss value. For example, given the prompt
One day, a little boy named Tim went to play with his friend, Sam. They wanted to play a game with a ball. The game was to see who could get the best score. Tim threw the ball and Sam tried to catch it. Sam missed the ball and it went far away. Sam was angry. He said, "I don’t like this game. I want to lie down and rest. "Tim said, "Okay, let’s lie down and rest. Then we can try again to get a good score." They lay down under a big tree and looked at the sky. They saw birds flying and clouds moving. After a while, they got up and a masked mixer yields a fairly coherent story completion:
@@
whereas a transformer with a slightly higher loss yields the the following completion with somewhat more deviation from the plot.
@@
4.3 Masked mixers are more efficient causal language modelers for small (<512 tokens) but not large-context training
To test the ability of masked mixers to be trained efficiently for causal language modeling on datasets that better represent those used for large frontier models while maintaining a reasonable training time on our limited compute we turned to a 10 billion token subset of the FineWeb-edu [24]. We train a Llama-3 [5] tokenizer of size 8000 on this dataset before using this tokenizer to train randomly-initialized transformers and mixers.
We first profiled the Llama transformer model to determine the optimal number of attention heads for this larger dataset, and found that, as for TinyStories, a 4-headed model achieved the lowest per-step loss and required the least time per step (Figure S7). We next trained transformers and masked mixers on the FineWeb dataset with varying context windows and batch sizes according to the rule $n_{ctx}*n_{b}=65536$ to prevent memory overflow, and the results of these experiments are shown in Figure 4. We observe that as the language model’s mapping goes from $\mathbb{N}^{1024}\to\mathbb{N}$ to $\mathbb{N}^{128}\to\mathbb{N}$ (and becomes closer to a bijective and thus invertible map) the masked mixer’s relative performance increases compared to the transformer.
Figure 4: FineWeb Causal Language Model Training.
4.4 Masked mixers are more efficient language learners than early transformer implementations
In some sense it is unsurprising that current transformer implementations are somewhat more efficient learners for large context windows than masked mixers because transformers have seen a number of architectural improvements and compute optimizations since their introduction. It is useful therefore to compare the learning efficiencies of early transformer implementations of all-next-token causal language modeling to masked mixers with the same training method, as these can be thought of as being similar in their developmental stage. We chose a model that was introduced before improvements such as Rotary Positional Encoding [26] and Flash Attention 2 and sophisticated parameter initialization techniques to name a few.
Specifically we test the Huggingface implementation of GPT-1 [2], a CLM-style transformer model which was introduced shortly after the original transformer architecture. We find that GPT lags well behind the masked mixer’s training and validation loss values both for default as well as learning rate, batch size, and head number-optimized versions of the transformer, for both small and large datasets (Figure 5, Table S6).
Figure 5: GPT-1 and masked mixer training efficiencies for $n_{ctx}=512$.
5 Alignment between Model and Task Invertibility Predicts Training Efficiency
5.1 Neither Dataset nor Task Stochasticity do not explain mixer transformer performance differences
One explanation of the improved performance of masked mixers on smaller context windows and the improved performance of transformers on larger context windows is that transformers are better able to filter out noise via attention, and this noise could exist in the dataset itself or else in the language modeling task. We first investigated whether the gaps between transformer and masked mixer causal language model training efficiency were due to stochasticity inherent in the dataset, by which we mean the somewhat random structure inherent in natural languages. For example, observe that many equally valid next words could be used for most of the words in most English sentences such that any one of these could be picked at random, as nouns and verbs and adjectives are often free interchanged and grammars are relatively inconsistent between passages.
On straightforward way to test this is by observing CLM loss for each model type during training on a dataset that has a higher proportion of formal language text with fewer degrees of freedom. We chose to train on mathematical text, the FineMath 4+ dataset with approximately 10 billion tokens [27], as an example of such a dataset. After CLM training we find that lower cross-entropy loss is achieved as expected for a more deterministic dataset, but that once again transformers are more efficient for longer-context but not shorter-context training (Figure S6), supporting the hypothesis that inherent dataset stochasticity does not contribute to the transformer-mixer training efficiency gap.
We next addressed question of whether the masked mixer and transformer training efficiency gaps were caused by the language modeling task itself, in this case using all left tokens to predict one next token. We did this by modifying the language modeling task to a two-sided prediction, rather than left-to-right prediction that is usually performed during CLM, while training on the Fineweb. Bidirectional prediction may be thought of as being more or less equivalent to the masked token prediction performed by encoding models such as BERT [28] except that only one token is ‘masked’ per input.
Rather than use a masked token prediction method, however, we note that the training for this objective function is inherently inefficient: only a small percentage (no more that 15% for BERT) of tokens are masked for a given sequence such that most tokens are not trained during each forward pass and merely provide context for the masked tokens. We introduce a method by which every input token is trained upon in one pass, which allows for $1/0.15=6.67$ times the training throughput as a standard masked language modeling approach.
The key to this method is to note that one can use the all-next-token approach of causal language modeling in both forward and reverse directions as long as there is no more than one linear combination of forward and reverse modules per model. This is because, without loss of generality, information for a ‘next’ token can travel between $t_{n+1}$ and $t_{n-1}$ and thence to $t_{n}$ if there are more than one linear combinations of forward and reverse modules (see Figure S8 for a depiction of this process in the case of the masked mixer). Thus we split our bidirectional models into two parts, each which perform next token prediction but one that does so in the forward and one in the reverse direction.
As shown in Figure 6, we see that there is an almost identical loss gap between masked mixer and transformer for bidirectional training compared to left-token CLM training, providing evidence for the idea that task stochasticity does not explain the relative performance of each model type. Together with the earlier results in this section, we conclude that neither dataset nor task-specific stochasticity explain the mixer versus transformer training efficiency differences.
Figure 6: Top: bidirectional modeling approach with respect to token indices. Bottom: Masked Mixer and Transformer FineWeb loss during training.
5.2 Multi-token and many-token prediction efficiencies
After observing that masked mixers perform worse relative to transformers as $n$ increases during CLM training in which models learn $\mathbb{N}^{n}\to\mathbb{N}^{1}$ mappings, we investigated the relative training efficiencies of these models on language modeling tasks described as multifunctions of $\mathbb{N}^{n}\to\mathbb{N}^{m},m>1$ and more specifically $\mathbb{N}^{512}\to\mathbb{N}^{2}$.
The assumptions behind training a language model to predict one next token at a time is that this is an easier task than predicting many next tokens at once, and that simply being able to predict each next token allows for effective inference. The latter implies that a greedy strategy during training is best, but the ability of inference methods such as beam search to outperform greedy inference suggests that the training a model to predict multiple ‘next’ tokens may be beneficial. This has recently been referred under various names such as ‘multiple token prediction’ or ‘non-myopic pretraining’, and there are many method by which one could implement this idea ranging from a multi-headed approach where each head predicts one token [29] to the more direct approach of performing $n$ forward passes and add the corresponding loss values for each of $n$ tokens predicted, performing label shifting with a shift position size corresponding to the number of forward passes similar to that in [30]. We take the latter approach and perform two forward passes through the model for two-token prediction. As shown in Figure 7 (a), the loss gap between masked mixer and transformer is somewhat smaller for two-token prediction compared to one-token prediction at 1.8% and 2.0%, respectively (refer to Figure 4).
We also investigated the efficiencies of language models trained to predict many next tokens in one forward pass instead of only one such that each token prediction is $\mathbb{N}^{n/2}\to\mathbb{N}$ with all token predictions being $\mathbb{N}^{n/2}\to\mathbb{N}^{n/2}$. Somewhat arbitrarily choosing to generate half of the text (meaning a context window of 512 we use the first 256 tokens as inputs and perform loss computation on the next 256 tokens), we find that a masked mixer and transformer yield nearly identical Fineweb training efficiencies Figure 7 (c).
Figure 7: FineWeb training curves for sequential multi-token prediction and parallel many-token prediction. Masked Mixers are $d_{m}=1024,n_{l}=16$ and transformers are llama-style $d_{m}=512,n_{l}=16$ with $n_{ctx}=512$ for both experiments.
5.3 Masked Mixers are far better Autoencoders than Transformers
At the other extreme of the $\mathbb{N}^{n}\to\mathbb{N}^{1}$ mapping that is performed during next token prediction, we test masked mixer and transformer training efficiencies on a task that requires an $\mathbb{N}^{n}\to\mathbb{N}^{n}$ (i.e. invertible) transformation: a non-trivial (with respect to inter-token information flow) autoencoding in which an encoder takes the input and compresses it to one hidden layer, and the decoder takes this hidden layer and attempts to regenerate the input tokens. Such an autoencoder may be implemented by using a encoder-decoder model where a standard masked mixer acts as the encoder, the last hidden state of the last token is taken as the embedding, and a decoder composed of a stack of masked mixer modules transforms this embedding (repeated for each input word) into a sequence to match the input as shown in 8. We implement an analogous architecture using a modified Llama-style transformer without word-token embedding layers or language modeling heads for encoder and decoder modules.
This mixer autoencoder reaches a cross-entropy loss on FineMath comparable to non-optimized transformers and mixers (albeit with around twice the number of training steps required) and very little or no overfitting, indicating that the embedding created by the model’s encoder provides generalizable information on the input to the decoder, and in particular enough information for the decoder to reconstruct the entire input sequence approximately as accurately as trained all-next-token causal language models. In contrast, transformers are far less efficient for this autoencoding and are unable to predict input tokens with anywhere near the accuracy of causal language modeling-trained transformers or masked mixers or mixer autoencoders (Figure 8). We observe this to be the case for TinyStories as well, and confirm that these differences are not caused by differences in $d_{m}$ between transformer and masked mixer (Figure S13).
Figure 8: Left: Mixer Autoencoder architecture. Right: Autoencoder loss on FineMath 4+.
6 Masked Mixers are Better for Retrieval
@@
On the other hand, matching a sequences of tokens to another sequence of tokens may be thought of as approximations to a bijective function which would not be expected to benefit from information reduction or non-invertibility. Masked mixers have been shown to be effectively invertible for small inputs and in that respect retain much more input information than mixers, suggesting that these models are better suited to retrieval than models with attention assuming that retrieval requires most input information. In this section we test this idea experimentally two ways, and find that indeed last-hidden-layer embeddings from masked mixers are far superior for retrieval tasks compared to the same embeddings from transformers in both settings.
@@ 6.1 Synthetic dataset generation and embedding
To investigate retrieval abilities of transformers and masked mixers on in-distribution data while maintaining flexibility as to the number of samples included per retrieval, we generated a synthetic summary sentence for the first 200k samples (truncated to 512 tokens maximum) in the Finewen dataset using Llama-3 (8b) Instruct [5] by prompting that model to read each 512-token-length text segment and give a one-sentence summary of that segment. We found that this relatively small model was more than capable of accurately and concisely summarizing each segment. We repeated this procedure with the FineMath 4+ dataset, for either 200k or 400k samples, and changed the model prompt to make the summary query more concise (and more difficult to match).
Each text segment and its corresponding summary were fed to a pretrained Llama-style transformer or masked mixer and the last hidden layer activations were saved as the model’s embedding. We use the second-to-last token’s output for the embedding because the ’last’ masked mixer token is untrained due to the index shift that occurs during all-next-token training of these models. This procedure only rarely would be expected to lead to input information loss, as the synthetic summary queries nearly all contain ending punctuation tokens. As masked mixers have an absolute positional encoding, we pretrain these models on left-padded tokens rather than the right-padded token segments used elsewhere in this work in order to be able to use the second-to-last token hidden layer for inputs of all sizes. We experimentally determined that left versus right padding during pretraining yields virtually no difference for subsequent retrieval training for transformers, as long as the retrieval training itself proceeds with left padding. For consistency, we pretrained transformers used for retrieval with left-padded inputs as well.
6.2 Indirect Retrieval Model Architecture and Training Approach
We first investigated the ability of masked mixers and transformers to perform retrieval after next-token causal language model pretraining where the model performing the embedding remains unchanged during retrieval training. This allows us to address the question of how effective each model’s pretrained embeddings are for retrieval while us to reference these values to existing invertibility measurements (Figure S5). We note that there are large training throughput gains with this method as well as retrieval training typically proceeds by modifying the parameters of the embedding model directly [32] which is efficient for inference but in some sense is inefficient for training because one forward pass is required for each input for each batch of matching and non-matching sequences. If one first embeds each input using a trained generative model and then trains a separate model on these embeddings then a single forward pass is required for comparison of all inputs in each forward pass of an input size $c=n_{ctx}$ where $c$ is the number of embeddings.
@@ 6.2 Indirect Retrieval Model Architecture and Training Approach
For our retrieval model we chose to implement a mixer with bidirectional convolutions rather than masked ones, similar to MLP mixers but with only one convolutional layer rather than two between sequence elements. This was chosen because a comparison of all possible matches requires information from all input elements by definition, making transformer models unsuitable. This was experimentally verified, as transformers were found to be very poor retrieval models even for small comparison contexts (Table S7). Word -token embedding layers are removed (as the inputs are embeddings already) as are language heads, and instead each last hidden layer has a $d_{model}\to 1$ transformation where the unitary output corresponds to the likelihood (after Softmax transformation) of that input embedding being the correct match. Softmax transformation serves to stabilize the retrieval model’s loss, as increasing all logit values does not decrease the cross-entropy loss, and a depiction of this architecture for masked mixers is found in Figure 9. We make use of the Pytorch Cross-Entropy Loss function’s compatibility with class probabilities for this model.
@@
The outputs after Softmax transformation are compared to the true distribution via standard cross-entropy loss, which is then back-propagated through the retrieval model but not further to the generative model or input embedding. Note that the logits are not Softmax-transformed during inference as this operation has no effect on top-k choice without noise introduction. We use a relatively modest batch size (512) compared to most retrieval model training approaches. In contrast to the rest of this work, the efficiency of an embedding for retrieval modeling is not measured in loss per compute amount or even loss per number of samples seen but instead minimum evaluation cross-entropy loss over a pre-determined.
The evaluation set cross-entropy loss after indirect retrieval training for a $d_{m}=512$ masked mixer and two Llama-style transformers pretrained on the FineWeb dataset are shown in Figure 9 and give evidence for the hypothesis that masked mixer embeddings are superior for retrieval training compared to those from transformers. Notably, training fails (with no training loss minimization) when using embeddings from models that were not pretrained, or if the retrieval model is implemented with a transformer architecture instead of a mixer.
Model $n=32$ 128 256 512 1024
Masked Mixer 0.28 0.58 0.78 0.88 1.20
Transformer, h=4 0.55 1.24 1.76 1.84 6.93
Transformer, h=32 0.61 1.19 1.59 1.88 6.93
Figure 9: Retrieval mixer architecture and FineWeb retrieval evaluation cross-entropy loss (sampled without replacement).
6.3 Masked Mixers are more suitable for cosine similarity InfoNCE -trained retrieval
The two major downsides to the secondary retrieval model training approach is that there is no guarantee that the embedding model is capable of capturing the information necessary to perform retrieval, as the embedding model was trained for an entirely different task (causal language modeling in this case), and that inference requires much more computation. Empirically the first downside is observed to be the case as the retrieval model tends to overfit the retrieval dataset. It is therefore pertinent to observe the ability of the transformer and masked mixer to be trained to form embedding suitable for retrieval and one particularly promising way to do this is to perform noise contrastive training on a causal language-pretrained model.
$\mathbb{L}=-\log\frac{f(q^{+},d^{+})}{f(q^{+},d^{+})+\sum_{i}f(q^{+},d_{i})}$ (7)
We implement a temperature-scaled cosine distance-based InfoNCE loss where $q^{+}$ indicates the query embedding, $\tau=0.02$ the temperature hyperparameter set to 0.02, $d^{+}$ the matching text embedding and $d_{i}$ randomly sampled nonmatching text embeddings (Equation (7)).
$f(a,b)=\exp\left(\frac{1}{\tau}\cos(O(a,\theta),O(b,\theta))\right)$ (8)
We use the model’s last hidden layer of the second-to-last token as the embedding as shown in Equation 8 instead of appending an extra token to each sequence, and sample with a context window size of 512. We include no hard negatives and sample one query with one matching and thirty non-matching text segments, and process four batches per optimizer update (one per GPU). We sample each batch using an analogous method to Algorithm 1 except that token sequences rather than embeddings are retrieved, and each input is assembled in the batch dimension such that a batch size of 32 has 32 separate text sequences.
We use a standard approach to retrieval inference via the cosine similarity identity defined in Equation (9) which we implement in an efficient batched form by first normalizing last hidden layer activations of the query $x$ and n targets $y_{n}$ such that $||x||=||y_{n}||=1$ before performing vector multiplication on the query output $x$ and row-concatenated target vectors $Y$ such that the output is $z=xY^{T}$, where the top-1 value of the output vector $z$ is found in the batch dimension.
$z_{n}=\cos(x,y_{n})=\frac{x\cdot y_{n}}{||x||\;||y_{n}||}$ (9)
We first benchmarked a state-of-the-art (when this work was commenced) large model pretrained on billions of tokens and postrained via cosine similarity InfoNCE for retrieval (among other tasks), e5 Mistral Instruct [32]. When we applied to the synthetic Fineweb retrieval dataset introduced in the last section we observed that top-1 evaluation accuracy for 32- and 64-sample retrieval of around 100% for this model (compared to 92.8% and 87.7% respectively for the masked mixer as trained using the method above) suggesting that this dataset is not particularly challenging for a strong model like e5 Mistral Instruct. We therefore generated a new synthetic dataset in which the same Llama model was instructed to generate shorter queries for FineMath 4+ text segments. This can be viewed as being a more difficult retrieval task because the dataset itself is restricted in subject and also because the queries are themselves less shorter and descriptive than they were for the FineWeb.
The Top-1 retrieval accuracy of each model with the FineMath 4+ evaluation Cross-Entropy Loss for models trained on this dataset are shown in Table 2. For both top-1 retrieval accuracy and loss values, measurements were performed on evaluation datasets held out during each respective training, with the pretraining corpus including both training and evaluation target text segments used for retrieval and none of the queries. All models except the already-trained e5 Mistral Instruct were trained using n=180k retrieval samples (one sample being a query with its matching target text) except where noted, and evaluated on a hold-out dataset of 20k samples (n=400k implies 380k training and 20k evaluation samples). Note that e5 Mistral Instruct uses a different tokenizer such that its cross-entropy loss is not comparable, and that due to the masked mixer’s far fewer inter-token parameters the transformer in the first line of Table 2 was pretrained using roughly the same compute as the $n_{l}=32$ masked mixer and $d_{m}=1024$ masked mixer and around double the compute as the masked mixer in the second line of that table.
Model Top-1 Retrieval Accuracy (%) Pretraining Loss
Transformer 70.4 1.39
Masked Mixer 81.2 1.79
Masked Mixer $n_{l}=32$ 84.6 1.65
Masked Mixer $d_{m}=1024$ 86.0 1.53
Transformer s=400k 95.0 1.39
e5 Mistral Instruct $d_{m}=4096$ 95.1
Masked Mixer $d_{m}=1024$, s=380k 98.2 1.53
Table 2: FineMath evaluation retrieval top-1 accuracy with n=32 sample size per retrieval, postrained on s=180k synthetic retrieval samples with $d_{m}=512$ except where denoted.
For practical query retrieval tasks one usually wants to perform retrieval using sample sizes of hundreds or thousands of targets. This is a major drawback for benchmark datasets which compare queries to only a few targets (for example as is typical for datasets like MSMarco [33]) because few-sample retrieval performance has no guarantee of extrapolating to many-sample retrieval. It is therefore important to measure many-sample retrieval performance directly, and the these results are shown in Tables 1 and 3. Most notably, the transformer model’s near-parity performance to the much larger e5 Mistral Instruct on small-context inputs vanishes as the sample size increases such that it becomes far worse for large-context retrieval relative to e5 Mistral, whereas the masked mixer (trained on identical data) exhibits uniformly superior performance to the other models across all sample sizes. As has been observed for transformers, increasing the masked mixer’s pretraining compute increases the resulting model’s retrieval performance (Table 3), and pretraining on an I.I.D. dataset to the retrieval task yields increased re trieval performance than pretraining on a dataset with a mix of in- and out of- distribution data such as that in the FineWeb (Table S7).
Model $n=32$ 64 128 256 512 1024 2048 4096 8192
Transformer 95.0 92.3 88.9 84.5 79.1 72.6 65.7 57.6 50.4
Transformer Autoencoder 94.4 92.0 88.4 84.3 79.1 73.1 67.0 60.6 53.3
Mixer Autoencoder 97.1 95.4 93.1 90.0 86.3 82.1 76.7 70.7 64.5
Mixer Autoencoder, pretrained 500k 96.9 95.5 93.1 90.0 86.6 82.0 76.8 71.5 65.0
e5 Mistral Instruct 95.1 93.3 90.9 88.5 85.6 82.2 78.4 73.5 68.8
Masked Mixer 98.2 97.2 95.8 93.4 90.2 86.5 81.7 76.6 70.6
Masked Mixer, pretrained 500k 98.2 97.3 95.8 93.5 90.7 86.8 82.4 77.3 71.7
Table 3: FineMath 4+ evaluation retrieval top-1 accuracy (%) for various sample sizes $n$, transformer and mixers were pretrained for 200k steps except where noted, and post-trained on s=380k synthetic retrieval samples
6.4 Autoencoders yield effective pretrained embeddings for InfoNCE
We also investigated the ability of a mixer pretrained on the FineWeb dataset using the autoencoding training presented in Section 5.3, modifying the InfoNCE training protocol accordingly to use the last token of the model’s encoder rather than second-to-last token’s embedding, and discarding the decoder. After InfoNCE retrieval training on FineMath, we find that this model achieves a higher 32-sample top-1 FineMath 4+ retrieval accuracy than the masked mixer pretrained via next token prediction (Table S7), despite the encoder containing half the number of layers as the CLM-pretrained masked mixer, and is even more accurate than the transformer trained on FineMath itself. Pretraining the mixer autoencoder on FineMath yields better accuracy than the transformer applied to the same dataset, but less accuracy than the CLM-pretrained masked mixer (Table 3).
Masked mixers and transformers without pretraining are virtually untrainable using the cosine similarity-based InfoNCE method, as these models fail to minimize the objective function value to any significant degree in our experiments. Curiously, transformers pretrained to autoencode using this method yield effective retrieval models after InfoNCE training and indeed are more accurate large-sample retrievers than CLM-pretrained transformers, despite the very poor performance of the transformer with respect to minimization of the autoencoding objective function (Table 3, Figure 8). In general it seems that minimization of the autoencoding objective function does not yield benefits for retrieval beyond a certain point, as continuing the mixer autoencoder pretraining to 500 thousands steps yields a much lower evaluation loss (CEL of 1.38) compared to the standard 200k steps (CEL of 2.18) but very little change in retrieval performance post-InfoNCE (Table 3).
We conclude that autoencoding provides a promising new direction for pretraining for retrieval models as it appears to be far less sensitive to the amount of compute used, although more work needs to be done to more fully determine the strengths and weaknesses of this approach as compared to causal language modeling-based pretraining.